# On the applicability of the cumulant expansion method for the calculation of transport properties in electron-phonon systems

### v2

@@ IV.3.1 General remarks
The CE method in the context of the Fröhlich model was previously examined in Ref. [16], but only at zero temperature. In the following, however, we turn our attention to the finite-temperature transport properties of the Fröhlich model. Before proceeding, let us note that the results of Ref. [16], among other things, demonstrate that the quasiparticle properties obtained using CE are significantly closer to the numerically exact benchmark (QMC [61] in their case) than those predicted by the Migdal approximation. The fact that this observation is fully consistent with the conclusions of Sec. IV.1.1 for the Peierls model and of Ref. [37] for the Holstein model suggests that some general, largely model-independent conclusions can be drawn regarding the advantages and limitations of the CE method. Consequently, as we will see, the conclusions reached for the Peierls model also hold for the Fröhlich model. For this reason, we will not provide a detailed analysis as in Sec. IV.1, but will instead summarize only the main results.
@@ IV.3.2 Numerical results
The calculations were performed using parameters representative of the conduction bands in three polar semiconductors: GaAs ($\alpha=0.072,\;\omega_{0}=36\,\mathrm{meV},\;m_{0}/m_{e}=0.067$), GaN ($\alpha=0.407,\;\omega_{0}=91.2\,\mathrm{meV},\;m_{0}/m_{e}=0.20$), and ZnO ($\alpha=1.022,\;\omega_{0}=72.3\,\mathrm{meV},\;m_{0}/m_{e}=0.275$), where $m_{0}/m_{e}$ denotes the ratio of the electron band mass to the free-electron mass. In these materials, the ratio of the fully renormalized electron mass $m^{\ast}$ (including all interaction effects) to the band mass satisfies $m_{0}/m^{\ast}\gtrsim 0.8$ [61]. In light of this, and based on the results presented in Sec. IV.1 and Refs. [37, 39] for the Peierls and Holstein models, we expect the self-consistent Migdal approximation to provide accurate mobility estimates over the parameter range considered in this work. This point is important, as numerically exact mobility data for the Fröhlich model are not available.
@@ A Quasiparticle properties for the Peierls model
[…]

### v3

@@ IV.2 Analytical Rationale for Employing CE
Lastly, these considerations allow us to construct Fig. 10, which illustrates regions in the $(\lambda,T)$ parameter space where the CE approach is expected to be reliable. To this end, we introduce a scalar quantity $S$ that quantifies how close the truncated result $n_{e}(r_{\mathrm{max}})$ is to its asymptotic value $n_{e}(r\to\infty)$,
$S(r_{\mathrm{max}})=\exp\!\left[-\frac{\Delta n_{e}(r_{\mathrm{max}})+\Delta^{2}n_{e}(r_{\mathrm{max}})}{\eta\,|n_{e}(r_{\mathrm{max}})|}\right].$ (96)
Here, $\Delta n_{e}(r_{\mathrm{max}})=|n_{e}(r_{\mathrm{max}})-n_{e}(r_{\mathrm{max}}-1)|$, $\Delta^{2}n_{e}(r_{\mathrm{max}})=|n_{e}(r_{\mathrm{max}})-2n_{e}(r_{\mathrm{max}}-1)+n_{e}(r_{\mathrm{max}}-2)|$, while $\eta$ is the parameter defining relative convergence tolerance which we take to be $\eta=0.2$ to best illustrate our results. We emphasize that the inclusion of both $\Delta n_{e}$ and $\Delta^{2}n_{e}$ in Eq. (96) is essential. While $\Delta n_{e}(r_{\mathrm{max}})$ alone might appear sufficient as a convergence criterion, this can be misleading in practice. In particular, for not too strong couplings one often finds $|n_{e}(r)-n_{e}(r-1)|\ll|n_{e}(r)-n_{e}(r-2)|$, reflecting the fact that, for odd $r$, certain contributions to $\mathcal{M}_{r}(k)$ are odd in $k$ and therefore vanish upon summation in Eq. (94). The inclusion of $\Delta^{2}n_{e}$ accounts for this effect and thereby provides a more reliable convergence criterion.
Figure 10: Illustration of the regions where CE is expected to be accurate in the Peierls model for $\omega_{0}=0.5$. $S$ is given by Eq. (96), with $r_{\mathrm{max}}=12$ and $\eta=0.2$.
@@ IV.3.1 General remarks
The CE method in the context of the Fröhlich model was previously examined in Ref. [23], but only at zero temperature. In the following, however, we turn our attention to the finite-temperature transport properties of the Fröhlich model. Before proceeding, let us note that the results of Ref. [23], among other things, demonstrate that the quasiparticle properties obtained using CE are significantly closer to the numerically exact benchmark (QMC [70] in their case) than those predicted by the one-shot Migdal approximation. The fact that this observation is fully consistent with the conclusions of Sec. IV.1.1 for the Peierls model and of Ref. [45] for the Holstein model suggests that some general, largely model-independent conclusions can be drawn regarding the advantages and limitations of the CE method. Consequently, as we will see, the conclusions reached for the Peierls model also hold for the Fröhlich model. For this reason, we will not provide a detailed analysis as in Sec. IV.1, but will instead summarize only the main results.
@@ IV.3.2 Numerical results
The calculations were performed using parameters representative of the conduction bands in three polar semiconductors: GaAs ($\alpha=0.072,\;\omega_{0}=36\,\mathrm{meV},\;m_{0}/m_{e}=0.067$), GaN ($\alpha=0.407,\;\omega_{0}=91.2\,\mathrm{meV},\;m_{0}/m_{e}=0.20$), and ZnO ($\alpha=1.022,\;\omega_{0}=72.3\,\mathrm{meV},\;m_{0}/m_{e}=0.275$), where $m_{0}/m_{e}$ denotes the ratio of the electron band mass to the free-electron mass. The effective mass parameters for GaAs and GaN were taken from Ref. [71] while other parameters for GaAs and GaN (energy of longitudinal optical phonon, static and high-frequency dielectric constant) were taken from Ref. [72]. The parameters for ZnO were taken from Ref. [73]. In these materials, the ratio of the fully renormalized electron mass $m^{\ast}$ (including all interaction effects) to the band mass satisfies $m_{0}/m^{\ast}\gtrsim 0.8$ [70]. In light of this, and based on the results presented in Sec. IV.1 and Refs. [45, 47] for the Peierls and Holstein models, we expect the self-consistent Migdal approximation to provide accurate mobility estimates over the parameter range considered in this work. This point is important, as entirely real-axis numerically exact mobility data for the Fröhlich model are not available. We note that, although quantum Monte Carlo (QMC) is a highly reliable method for calculations on the imaginary frequency axis [74], the use of analytical continuation can severely compromise the accuracy of the resulting dc mobility (see, e.g., Fig. 15 of Ref. [14] and Sec. II C of the Supplemental Material of Ref. [40]). This effect is clearly illustrated in the Holstein model at weak–intermediate coupling, $\lambda=1/2$: Fig. 1 of Ref. [41] shows that the hierarchical equations of motion (HEOM) method is in excellent agreement with quantum typicality—another numerically exact approach that does not rely on analytical continuation—whereas Fig. S5 of the Supplemental Material of Ref. [46] demonstrates that QMC underestimates the mobility in the same regime by nearly an order of magnitude. Regardless, the QMC results of Ref. [74] are not suitable as benchmarks here, as they were reported only for $\alpha=2.5$ and $6$, well outside the weak-to-intermediate-coupling regimes relevant for assessing the accuracy of the CE approach.
@@ A Quasiparticle properties for the Peierls model
In Sec. IV A 1 of the main text, we showed the cumulant expansion (CE), self-consistent Migdal approximation (SCMA), Migdal approximation (MA), and generalized Green’s function cluster expansion (GGCE) benchmark predictions for the band-mass to renormalized-mass ratio $m_{0}/m^{*}$ (note that $m_{0}=1/(2J)$) and the ground-state energy $E_{p}$ for $\omega_{0}=0.5$ and $T=0$. Here, in Figs. S4– S4 we additionally show results for $\omega_{0}=1.0$, $\omega_{0}=0.044$, and $\omega_{0}=3.0$. Also, in Fig. S4 we illustrate CE lifetime for $J=g=1$, $T=2$; see Eq. (63) from the main text.